# Computing an Aircraft’s Gliding Range and Minimal Return Altitude in Presence of Obstacles and Wind

### v1

@@ II.4 Minimal return altitude problem
To give some intuition about this equation let us suppose the wind is zero, and therefore $g$ is constant. Suppose also that there are no obstacles ($h_{\text{min}}=0$). In this case the solution is $V(\bm{y})=\frac{1}{g}\left\lVert\bm{y}-\bm{x}_{a}\right\rVert$, that is, a cone of slope $1/g$ with the vertex at $\bm{x}_{a}$. When obstacles are present, the cone solution can fall below $h_{\text{min}}$, in this case it’s not valid anymore and $V$ instead starts tracking the profile of $h_{\text{min}}$. See figure 6 for a sketch of $V$ in the one dimensional case. One can see the process of solving MRAP as a front propagating outwards from $\bm{x}_{a}$, where the arrival time of the front is the minimal return altitude. Similarly to GRRP, in MRAP one defines the vector field $\bm{\hat{a}}_{M}:\operatorname{\mathbb{R}}^{2}\mapsto S^{1}$, where $\bm{\hat{a}}_{M}(\bm{y})$ is the vector that attains the maximum in (18). Integrating $-\bm{\hat{a}}_{M}(\bm{y})$ from the aircraft position, gives a feasible path terminating in $\bm{x}_{a}$. Feasible means that in all points the path does not go below $h_{\text{min}}$ and that the steepness is at least $1/g$. This path is however different in general from the path that minimizes the altitude loss to reach $\bm{x}_{a}$. In appendix H we compare the trajectories found in GRRP with those of MRAP and comment on their differences.
@@ III Algorithms
Equations (13) and (18) can be solved analytically only in the case of uniform wind and no obstacles. We present their solution in Appendix B. In all other cases, a numerical approach is needed. Algorithms to solve the HJB equations of the form (13) in absence of constraints have been developed by Sethian and Vladimirsky in [1, 2, 27]. These methods are based on discretizing the 2d plane using a grid of points and subsequently approximating the value of $U$ at each grid point. The algorithms mimic the front propagation by setting the values of $U$ starting from the nodes closer to the initial point (respectively the aircraft position and the airfield position in MRAP and GRRP) and expanding outwards.
@@
Glikonal-M, the algorithm to solve MRAP, is only able to deal with the zero wind case. This is due to the fact that solving the equations with an arbitrary wind field presents some additional technical challenges. In Appendix A.3 we derive the simplified form of the HJB equations (18) for the windless case. Glikonal-M takes as input the position of the airfield, the elevation profile and the glide ratio $g\in\operatorname{\mathbb{R}}$ (which is constant in absence of wind). It outputs an approximation of $V$ solving (18),(19) on a grid. The pseudocode of Glikonal-M is reported in Appendix E.
The grid spacing is a a parameter of both algorithms. The finer the grid, the smaller the error, the longer it takes to run. When the grid spacing goes to zero both algorithms exactly solve the respective PDEs. For a proof of this fact see [1, 2]. The time complexity of both algorithms is $O(n\log n)$, where $n$ is the number of points in the grid.
@@ IV Numerical results
In this section we present the numerical results obtained with Glikonal-G and Glikonal-M in several settings. The code to reproduce all experiments can be found at https://github.com/giovannipiccioli/glikonal.
@@ IV.1 Glikonal-G
In the first set of numerical experiments we assume for simplicity (and for visualization purposes) that the glide ratio is $1$ in absence of wind and that the aircraft’s airspeed is also $1$. When the wind is present we assume that the aircraft travels in every direction at fixed airspeed, therefore $g(\bm{y},z,\bm{\hat{a}})=\left\lVert\bm{\hat{a}}+\bm{W}(\bm{y},z)\right\rVert$. Notice that this is not the optimal glide ratio, as in principle, to maximize the glide ratio, one should increase the airspeed when moving upwind.  [^2] We choose not to optimize the glide ratio with respect to the airspeed since this would force us to specify the functional form of the glide polar $s:\operatorname{\mathbb{R}}\mapsto(-\infty,0)$ of the aircraft. Moreover it is reasonable to assume that a pilot would try to maintain a constant airspeed. In the following we will refer to $U$ as the solution of GRRP, and to $U_{G}$ as its approximation returned by Glikonal-G. We run our experiments on a $101\times 101$ grid with horizontal and vertical spacings of $1$. In all experiments the initial altitude is taken to be $z_{0}=100$.
@@
Panel (a) depicts the values of $z_{0}-U_{G}$, and its contour lines. The red dot is the initial position of the glider. As expected, the altitude decreases faster when travelling upwind, than downwind. The turquoise lines represent the optimal trajectories, and the red zone is the region that cannot be reached in gliding flight. In absence of obstacles and with constant wind, the optimal trajectories are straight as expected under the analytic solution. Panel (b) shows the relative error $(U_{G}-U)/U$ of the algorithmic approximation. We observe that $U_{G}$’s relative error is always below 3%. Moreover the error is always positive, implying that $U_{G}>U$, and therefore the algorithms errs on the conservative side, overestimating the lost altitude. In Appendix G the total and relative errors of the algorithm are computed in a number of settings. In all the cases we consider, the error is below 4%, and it is always on the conservative side. Finally in panel (c) the 3d ’gliding cone’, i.e., the function $z_{0}-U_{G}$ is plotted. One can again observe how the cone is steeper in the upwind direction.
Figure 2: Glikonal-G’s solution in the case of non uniform wind and single mountain peak. The red dot represents the glider’s initial position. (a) Wind direction (constant with altitude) plotted on top of the elevation profile (displayed as a heatmap). (b) Contour lines of the function $z_{0}-U_{G}$ (in white) and optimal trajectories (in turquoise). The heatmap is the elevation profile, and the red shaded regions are those not reachable in gliding flight. (c) 3d representation of the function $z_{0}-U_{G}$, in orange and of $h_{\text{min}}$.
@@ IV.1.1 GRRP on real data
To conclude our study of the Glikonal-G algorithm, we present a real-world example, with elevation data taken from [28]. We consider a glider with a glide ratio of 20 in absence of wind, moving at a fixed airspeed of 100km/h. For this experiment we use a $295\times 305$ grid with a spacing of $150$ m. The wind, represented with white arrows in figure 3 (a), is uniform and forms an angle of 30° with the south direction; its strength is $15$ km/h below $750$ m, then increases linearly to $40$ km/h at $2000$ m, and stays constant at higher altitudes. In the same panel the elevation profile is shown. We require the glider to be at more than 200m above the ground in every point of the trajectory. The initial position of the glider is (46.59540N, 6.40050E) at an altitude of $z_{0}=2070$ m (red dot). The purple dot represents the airfield of Montricher, the closest one to the glider. The airfield is located on the opposite side of the Jura mountain range. Panel (b) depicts the contour lines of the function $z_{0}-U_{G}$ in black, and the optimal trajectories in blue. To reach the airfield, one must first pass the mountain range to the south and then turn sharply left. Since the straight path from the glider to the airfield is obstructed, we remark that any line-of-sight algorithm would fail to recognize that the airfield is within gliding range. Panel (c) shows a three dimensional visualization of $z_{0}-U_{G}$ and of the elevation profile.
@@ IV.1.2 Computational efficiency
Figure 3: Glikonal-G on a real-world example. Horizontal distance is measured in kilometers, while the altitude in meters (see colorbar). The red and purple dots represent respectively the glider’s and airfield’s positions. (a) Heatmap of the elevation profile of the Jura mountain range. The glider and the closest airfield are located on opposite sides of the range. The white arrows represent the uniform wind across the map. (b) Contour lines of the altitude function $z_{0}-U_{G}$ in black and optimal trajectories in blue. The red shaded region is unreachable. (c) 3d representation of the altitude of the glider ($z_{0}-U_{G}$) in orange, on top of the elevation profile. for visualization purposes the vertical axis is rescaled by 8 in comparison to the horizontal axes. The purple vertical line highlights the position of the airfield.
@@ IV.2 Glikonal-M
To test the Glikonal-M algorithm we use it to approximate the function $V$, defined in (17). We name $V_{G}$ the approximation of $V$ returned by the algorithm. We run our experiments on a $101\times 101$ grid with horizontal and vertical spacings of $1$. We start with the case of a flat terrain ($h_{\text{min}}=0$) and glide ratio equal to one. The top row in figure 4 depicts the results of this experiment. In (a) the contour lines and heatmap of $V_{G}$ are plotted, together with the feasible re-entry trajectories, which in this case extend radially from the airfield. Using the results in Appendix B we compute the analytic solution $V$ and use it to plot the relative error $(V_{G}-V)/V$ in panel (b). The error is below 4% in the whole space, moreover $V_{G}$ is always greater than $V$, thus erring of the safe side. In Appendix G the error is computed for several different settings in which the analytical solution is available. In all the cases the error is below 5% and always on the conservative side. Panel (c) concludes this example by showing the function $V_{G}$, which takes the shape of a cone having the tip at the airfield. A glider above the cone will be able to reach the airfield, one below will not.
@@ IV.2.1 MRAP on real data
In this experiment we apply the Glikonal-M algorithm to study the minimal altitude needed to reach the airfield of Montricher in Switzerland, represented by a purple dot in figure 5. The elevation profile of this region is given as a heatmap in figure 3 (a). The size of the grid and the spacing are the same as in the Glikonal-G case. $h_{\text{min}}$ is set to the terrain elevation plus $200$ m, and the glide ratio is 20. In (a) the contour lines are plotted in black, and the re-entry trajectories in blue. There are no obstacles to the south and west of the airfield, hence the contour lines are circle arcs and the re-entry paths are just straight lines. To the north west the minimal return altitude is increased due to the presence of mountains, which one must pass to reach the airfield. Comparing the feasible paths in (a) with the elevation profile in 3 (a) we remark that the trajectories found by solving MRAP avoid passing on the points with highest elevation, and instead route the glider through the mountain passes.
Figure 5: Glikonal-M on a real-world example. The horizontal distances are measured in kilometers. The purple dot represent airfield’s position. (a) Contour lines of the minimal altitude function $V_{G}$ in black and feasible trajectories in blue. (b) 3d representation of the function $V_{G}$ in orange, on top of the elevation profile. for visualization purposes the vertical axis is rescaled by 8 in comparison to the horizontal axes. The purple vertical line highlights the position of the airfield.
@@ V Conclusions
In this work we introduced the GRRP and MRAP problems and presented the Glikonal-G and Glikonal-M algorithms to solve them. These respectively compute the GRR from a given position and altitude, and the minimal altitude to glide to a given airfield. First the aircraft dynamics is modeled by assuming that discontinuous changes in ground speed are allowed. Then an optimal control formulation of the aforementioned problems is presented. The HJB PDEs are derived, and then solved in a discretized form using Glikonal-G and Glikonal-M. The precision of the solution is verified in several settings, and results are presented both in the case of artificial and real-world data.
Glikonal-G is fast enough to run in real time on an on-board computer, and display the reachable region to the pilot throughout the flight. Moreover it is more advanced than other algorithms on several fronts. First it surpasses line-of-sight methods, which only consider straight gliding paths (at most with an initial turn to the desired heading). In several situations (see figures 3, 8, 2) the GRR found with these methods is incomplete. The optimal trajectories in Glikonal-G can instead turn around obstacles, therefore outputting the correct GRR. Second, it can handle any wind vector field, by having the glide ratio depend arbitrarily on the position, altitude and course of the glider.
@@ Acknowledgements.
I am grateful to the people of Montricher’s airfield for inspiring this work. I would also like to thank Lenka Zdeborová for giving me the freedom to work on this project.
@@ A.1 MRAP from GRRP
where $h_{f}$ here is a free parameter and represents the altitude loss to reach the airport. MRAP basically consists of computing $V$ in the whole space. Intuitively, the further we are from $\bm{x}_{a}$, the larger $V$. However this formulation cannot be readily translated into an algorithm that computes the $V$, since we’re constraining the final point of the dynamics instead of the initial one. For this reason we now present a sequence of equivalent problems with the aim of translating (21) into a problem with a convenient algorithmic solution.
@@
with the usual definition $\bm{\hat{a}}=\bm{a}/\left\lVert\bm{a}\right\rVert$ and $h$ being the lost altitude. This dynamics is identical to (6) except for the addition of the factor $\lambda$ which allows the glide ratio to be smaller that $\tilde{g}$. This means that the aircraft can descend with glide ratio at most $\tilde{g}$, but can also descend more steeply.
@@
The set of possible paths allowed by the relaxed dynamics (23) contains the set of paths allowed under the (6). Therefore $V_{2}\leq V$. We now prove that $V\leq V_{2}$. The idea is very simple: given a path that is steeper that $1/g$ is some section, we can always replace it with a path that has the same horizontal projection, (i.e. passes through the same points in the horizontal plane) but maintains a higher altitude than the original path by having steepness exactly $1/g$. If the original path satisfied the constraints, so does the modified path, since it passes higher than the first. Then taking the minimum over paths with the dynamics (23) is equivalent to taking the minimum over paths with the dynamics (6).
We now formalize the above. Consider a path $\bm{x}:\operatorname{\mathbb{R}}_{+}\mapsto\operatorname{\mathbb{R}}^{2}$ satisfying (25). We now reparametrize $\bm{x}$ in order to satisfy the non relaxed dynamics. Define $\bar{h}$ as the minimal altitude the aircraft can lose when following the path $\bm{x}$. $\bar{h}$ obeys the equation
@@
and thus $\bar{\bm{x}}$ satisfies the dynamics (6). At this point we remark that $\bar{h}(h)\leq h$. To see this suppose that for a certain $h$, $\bar{h}(h)=h$ (this happens for example when $h=\bar{h}=0$). Then deriving $\bar{h}$ we get
@@
Now that the equality between $V$ and $V_{2}$ has been established, we shall introduce the time reversed problem. This dynamics starts at an altitude $\bar{z}_{0}$ and climbs with steepness at least $g$. $h$ here represents the gained altitude from the initial point.
@@
    where $\bm{a}^{\prime}=-\bm{a}$. In the second passage we used (23), in the third we used that $z_{0}=\bar{z}_{0}+h_{f}$. In the final passage we obtained (29) by taking into account that the vector $\bm{a}$ should be reversed since we’re travelling in the opposite direction.

### v2

@@ II.4 Minimal return altitude problem
To give some intuition about this equation let us suppose the wind is zero, and therefore $g$ is constant. Suppose also that there are no obstacles ($h_{\text{min}}=0$). In this case the solution is $V(\bm{y})=\frac{1}{g}\left\lVert\bm{y}-\bm{x}_{a}\right\rVert$, that is, a cone of slope $1/g$ with the vertex at $\bm{x}_{a}$. When obstacles are present, the cone solution can fall below $h_{\text{min}}$, in this case, it’s not valid anymore and $V$ instead starts tracking the profile of $h_{\text{min}}$. See figure 6 for a sketch of $V$ in the one-dimensional case. One can see the process of solving MRAP as a front propagating outwards from $\bm{x}_{a}$, where the arrival time of the front is the minimal return altitude. Similarly to GRRP, in MRAP one defines the vector field $\bm{\hat{a}}_{M}:\operatorname{\mathbb{R}}^{2}\mapsto S^{1}$, where $\bm{\hat{a}}_{M}(\bm{y})$ is the vector that attains the maximum in (18). Integrating $-\bm{\hat{a}}_{M}(\bm{y})$ from the aircraft position, gives a feasible path terminating in $\bm{x}_{a}$. Feasible means that in all points the path does not go below $h_{\text{min}}$ and that the steepness is at least $1/g$. This path is however different in general from the path that minimizes the altitude loss to reach $\bm{x}_{a}$. In appendix H we compare the trajectories found in GRRP with those of MRAP and comment on their differences.
@@ III Algorithms
Equations (13) and (18) can be solved analytically only in the case of uniform wind and no obstacles. We present their solution in Appendix B. In all other cases, a numerical approach is needed. Algorithms to solve the HJB equations of the form (13) in the absence of constraints have been developed by Sethian and Vladimirsky in [1, 2, 27]. These methods are based on discretizing the 2d plane using a grid of points and subsequently approximating the value of $U$ at each grid point. The algorithms mimic the front propagation by setting the values of $U$ starting from the nodes closer to the initial point (respectively the aircraft position and the airfield position in MRAP and GRRP) and expanding outwards.
@@
Glikonal-M, the algorithm to solve MRAP, is only able to deal with the zero wind case. This is because solving the equations with an arbitrary wind field presents some additional technical challenges. In Appendix A.3 we derive the simplified form of the HJB equations (18) for the windless case. Glikonal-M takes as input the position of the airfield, the elevation profile, and the glide ratio $g\in\operatorname{\mathbb{R}}$ (which is constant in the absence of wind). It outputs an approximation of $V$ solving (18),(19) on a grid. The pseudocode of Glikonal-M is reported in Appendix E.
The grid spacing is a parameter of both algorithms. The finer the grid, the smaller the error, the longer it takes to run. When the grid spacing goes to zero both algorithms exactly solve the respective PDEs. For a proof of this fact see [1, 2]. The time complexity of both algorithms is $O(n\log n)$, where $n$ is the number of points in the grid.
@@ IV Numerical results
In this section, we present the numerical results obtained with Glikonal-G and Glikonal-M in several settings. The code to reproduce all experiments can be found at https://github.com/giovannipiccioli/glikonal.
@@ IV.1 Glikonal-G
In the first set of numerical experiments, we assume for simplicity (and for visualization purposes) that the glide ratio is $1$ in the absence of wind and that the aircraft’s airspeed is also $1$. When the wind is present we assume that the aircraft travels in every direction at fixed airspeed, therefore $g(\bm{y},z,\bm{\hat{a}})=\left\lVert\bm{\hat{a}}+\bm{W}(\bm{y},z)\right\rVert$. Notice that this is not the optimal glide ratio, as in principle, to maximize the glide ratio, one should increase the airspeed when moving upwind.  [^2] We choose not to optimize the glide ratio with respect to the airspeed since this would force us to specify the functional form of the glide polar $s:\operatorname{\mathbb{R}}\mapsto(-\infty,0)$ of the aircraft. Moreover, it is reasonable to assume that a pilot would try to maintain a constant airspeed. In the following we will refer to $U$ as the solution of GRRP, and to $U_{G}$ as its approximation returned by Glikonal-G. We run our experiments on a $101\times 101$ grid with horizontal and vertical spacings of $1$. In all experiments, the initial altitude is taken to be $z_{0}=100$.
@@
Panel (a) depicts the values of $z_{0}-U_{G}$, and its contour lines. The red dot is the initial position of the glider. As expected, the altitude decreases faster when travelling upwind, than downwind. The turquoise lines represent the optimal trajectories, and the red zone is the region that cannot be reached in gliding flight. In the absence of obstacles and with constant wind, the optimal trajectories are straight as expected under the analytic solution. Panel (b) shows the relative error $(U_{G}-U)/U$ of the algorithmic approximation. We observe that $U_{G}$’s relative error is always below 3%. Moreover the error is always positive, implying that $U_{G}>U$, and therefore the algorithm errs on the conservative side, overestimating the lost altitude. In Appendix G the total and relative errors of the algorithm are computed in a number of settings. In all the cases we consider, the error is below 4%, and it is always on the conservative side. Finally in panel (c) the 3d ’gliding cone’, i.e., the function $z_{0}-U_{G}$ is plotted. One can again observe how the cone is steeper in the upwind direction.
Figure 2: Glikonal-G’s solution in the case of non-uniform wind and single mountain peak. The red dot represents the glider’s initial position. (a) Wind direction (constant with altitude) plotted on top of the elevation profile (displayed as a heatmap). (b) Contour lines of the function $z_{0}-U_{G}$ (in white) and optimal trajectories (in turquoise). The heatmap is the elevation profile, and the red-shaded regions are those not reachable in gliding flight. (c) 3d representation of the function $z_{0}-U_{G}$, in orange and of $h_{\text{min}}$.
@@ IV.1.1 GRRP on real data
To conclude our study of the Glikonal-G algorithm, we present a real-world example, with elevation data taken from [28]. We consider a glider with a glide ratio of 20 in the absence of wind, moving at a fixed airspeed of 100km/h. For this experiment, we use a $295\times 305$ grid with a spacing of $150$ m. The wind, represented with white arrows in figure 3 (a), is uniform and forms an angle of 30° with the south direction; its strength is $15$ km/h below $750$ m, then increases linearly to $40$ km/h at $2000$ m, and stays constant at higher altitudes. In the same panel, the elevation profile is shown. We require the glider to be at more than 200m above the ground in every point of the trajectory. The initial position of the glider is (46.59540N, 6.40050E) at an altitude of $z_{0}=2070$ m (red dot). The purple dot represents the airfield of Montricher, the closest one to the glider. The airfield is located on the opposite side of the Jura mountain range. Panel (b) depicts the contour lines of the function $z_{0}-U_{G}$ in black, and the optimal trajectories in blue. To reach the airfield, one must first pass the mountain range to the south and then turn sharply left. Since the straight path from the glider to the airfield is obstructed, we remark that any line-of-sight algorithm would fail to recognize that the airfield is within gliding range. Panel (c) shows a three dimensional visualization of $z_{0}-U_{G}$ and of the elevation profile.
@@ IV.1.2 Computational efficiency
Figure 3: Glikonal-G on a real-world example. Horizontal distance is measured in kilometers, while the altitude is in meters (see colorbar). The red and purple dots represent respectively the glider’s and airfield’s positions. (a) Heatmap of the elevation profile of the Jura mountain range. The glider and the closest airfield are located on opposite sides of the range. The white arrows represent the uniform wind across the map. (b) Contour lines of the altitude function $z_{0}-U_{G}$ in black and optimal trajectories in blue. The red-shaded region is unreachable. Background map from [29]. (c) 3d representation of the altitude of the glider ($z_{0}-U_{G}$) in orange, on top of the elevation profile. for visualization purposes, the vertical axis is rescaled by 8 in comparison to the horizontal axes. The purple vertical line highlights the position of the airfield.
@@ IV.2 Glikonal-M
To test the Glikonal-M algorithm we use it to approximate the function $V$, defined in (17). We name $V_{G}$ the approximation of $V$ returned by the algorithm. We run our experiments on a $101\times 101$ grid with horizontal and vertical spacings of $1$. We start with the case of a flat terrain ($h_{\text{min}}=0$) and glide ratio equal to one. The top row in figure 4 depicts the results of this experiment. In (a) the contour lines and heatmap of $V_{G}$ are plotted, together with the feasible re-entry trajectories, which in this case extend radially from the airfield. Using the results in Appendix B we compute the analytic solution $V$ and use it to plot the relative error $(V_{G}-V)/V$ in panel (b). The error is below 4% in the whole space, moreover, $V_{G}$ is always greater than $V$, thus erring on the safe side. In Appendix G the error is computed for several different settings in which the analytical solution is available. In all the cases the error is below 5% and always on the conservative side. Panel (c) concludes this example by showing the function $V_{G}$, which takes the shape of a cone having the tip at the airfield. A glider above the cone will be able to reach the airfield, one below will not.
@@ IV.2.1 MRAP on real data
In this experiment, we apply the Glikonal-M algorithm to study the minimal altitude needed to reach the airfield of Montricher in Switzerland, represented by a purple dot in figure 5. The elevation profile of this region is given as a heatmap in figure 3 (a). The size of the grid and the spacing are the same as in the Glikonal-G case. $h_{\text{min}}$ is set to the terrain elevation plus $200$ m, and the glide ratio is 20. In (a) the contour lines are plotted in black, and the re-entry trajectories in blue. There are no obstacles to the south and west of the airfield, hence the contour lines are circle arcs and the re-entry paths are just straight lines. To the northwest, the minimal return altitude is increased due to the presence of mountains, which one must pass to reach the airfield. Comparing the feasible paths in (a) with the elevation profile in 3 (a) we remark that the trajectories found by solving MRAP avoid passing on the points with highest elevation, and instead route the glider through the mountain passes.
Figure 5: Glikonal-M on a real-world example. The horizontal distances are measured in kilometers. The purple dot represents the airfield’s position. (a) Contour lines of the minimal altitude function $V_{G}$ in black and feasible trajectories in blue. (b) 3d representation of the function $V_{G}$ in orange, on top of the elevation profile. For visualization purposes, the vertical axis is rescaled by 8 in comparison to the horizontal axes. The purple vertical line highlights the position of the airfield.
@@ V Conclusions
In this work, we introduced the GRRP and MRAP problems and presented the Glikonal-G and Glikonal-M algorithms to solve them. These respectively compute the GRR from a given position and altitude, and the minimal altitude to glide to a given airfield. First, the aircraft dynamics is modeled by assuming that discontinuous changes in ground speed are allowed. Then an optimal control formulation of the aforementioned problems is presented. The HJB PDEs are derived, and then solved in a discretized form using Glikonal-G and Glikonal-M. The precision of the solution is verified in several settings, and results are presented both in the case of artificial and real-world data.
Glikonal-G is fast enough to run in real time on an on-board computer and display the reachable region to the pilot throughout the flight. Moreover, it is more advanced than other algorithms on several fronts. First, it surpasses line-of-sight methods, which only consider straight gliding paths (at most with an initial turn to the desired heading). In several situations (see figures 3, 8, 2) the GRR found with these methods is incomplete. The optimal trajectories in Glikonal-G can instead turn around obstacles, therefore outputting the correct GRR. Second, it can handle any wind vector field, by having the glide ratio depend arbitrarily on the position, altitude, and course of the glider.
@@ Acknowledgements.
I am grateful to the people of Montricher’s airfield for inspiring this work. I would also like to thank Lenka Zdeborová for granting me the freedom to work on this project.
@@ A.1 MRAP from GRRP
where $h_{f}$ here is a free parameter and represents the altitude loss to reach the airport. MRAP basically consists of computing $V$ in the whole space. Intuitively, the further we are from $\bm{x}_{a}$, the larger $V$. However, this formulation cannot be readily translated into an algorithm that computes the $V$, since we’re constraining the final point of the dynamics instead of the initial one. For this reason, we now present a sequence of equivalent problems with the aim of translating (21) into a problem with a convenient algorithmic solution.
@@
with the usual definition $\bm{\hat{a}}=\bm{a}/\left\lVert\bm{a}\right\rVert$ and $h$ being the lost altitude. This dynamics is identical to (6) except for the addition of the factor $\lambda$ which allows the glide ratio to be smaller than $\tilde{g}$. This means that the aircraft can descend with glide ratio at most $\tilde{g}$, but can also descend more steeply.
@@
The set of possible paths allowed by the relaxed dynamics (23) contains the set of paths allowed under the (6). Therefore $V_{2}\leq V$. We now prove that $V\leq V_{2}$. The idea is very simple: given a path that is steeper than $1/g$ is some section, we can always replace it with a path that has the same horizontal projection, (i.e. passes through the same points in the horizontal plane) but maintains a higher altitude than the original path by having steepness exactly $1/g$. If the original path satisfies the constraints, so does the modified path, since it passes higher than the first. Then taking the minimum over paths with the dynamics (23) is equivalent to taking the minimum over paths with the dynamics (6).
We now formalize the above. Consider a path $\bm{x}:\operatorname{\mathbb{R}}_{+}\mapsto\operatorname{\mathbb{R}}^{2}$ satisfying (25). We now reparametrize $\bm{x}$ in order to satisfy the non-relaxed dynamics. Define $\bar{h}$ as the minimal altitude the aircraft can lose when following the path $\bm{x}$. $\bar{h}$ obeys the equation
@@
and thus $\bar{\bm{x}}$ satisfies the dynamics (6). At this point, we remark that $\bar{h}(h)\leq h$. To see this suppose that for a certain $h$, $\bar{h}(h)=h$ (this happens for example when $h=\bar{h}=0$). Then deriving $\bar{h}$ we get
@@
Now that the equality between $V$ and $V_{2}$ has been established, we shall introduce the time-reversed problem. This dynamics starts at an altitude $\bar{z}_{0}$ and climbs with steepness at least $g$. $h$ here represents the gained altitude from the initial point.
@@
    where $\bm{a}^{\prime}=-\bm{a}$. In the second passage we used (23), and in the third we used that $z_{0}=\bar{z}_{0}+h_{f}$. In the final passage, we obtained (29) by taking into account that the vector $\bm{a}$ should be reversed since we’re travelling in the opposite direction.